OpenStudy (anonymous):

why we use g=9.81m/s^s near area of earth?

OpenStudy (anonymous):

The following video at 3:00 begins the derivation. I hope it helps. I can' t really explain it. http://www.khanacademy.org/science/physics/mechanics/v/mass-and-weight-clarification

OpenStudy (shane_b):

\[\large F_g=G\frac{m_{earth}}{r_{earth}^2}\]\[\large G (\text{gravitational constant})=6.673*10^{-11} m^3 kg^{-1}s^{-2}\]\[\large m_{earth}=5.9742x10^{24}kg\]\[\large r_{earth}=6378.1km\]\[\large F_g=(6.673*10^{-11} m^3 kg^{-1}s^{-2})\frac{5.9742x10^{24}kg}{(6,378,100m)^2}=\text{~}9.8m/s^2\]

OpenStudy (ghazi):

@shane_B if i am not mistaken, i guess unit of Force is N , secondly \[F=G \frac{ m1*m2 }{ r^2 }\] force of attraction between two heavenly bodies

OpenStudy (ghazi):

\[a= \frac{ F }{ m1 }= G \frac{ M _{earth} }{ R^2 _{earth} }\]

OpenStudy (ghazi):

\[M _{2}=M _{earth}\]

OpenStudy (ghazi):

@vannayen

OpenStudy (shane_b):

@ghazi: You're right...I oversimplified it and should have included the second mass in the calculation. I basically assumed a mass of 1kg for m2 and left it out of the equation. The answer should have been in kg*m/s^2 (N) which when divided by 1kg = 9.8m/s^2. @demitris: Nice image and good point about the radius varying between the equator and the poles. @vannayen: Note that this is not how we actually *measure* the force of gravity. That can be done using several other methods (atwood machine, springs, etc).